### v2

@@ Abstract
We introduce and formalize the notion of resource augmentation for maximin share allocations — an idea that can be traced back to the seminal work of Budish [JPE 2011]. Specifically, given a fair division instance with $m$ goods and $n$ agents, we ask how many copies of the goods should be added in order to guarantee that each agent receives at least their original maximin share, or an approximation thereof. We establish a tight bound of $m/e$ copies for arbitrary monotone valuations. For additive valuations, we show that at most $\min\{n-2,\left\lfloor\frac{m}{3}\right\rfloor(1+o(1))\}$ copies suffice. For approximate-MMS in instances with additive valuations, we give a tradeoff between the number of copies needed and the approximation guarantee. In particular, we prove that $\lfloor n/2\rfloor$ copies suffice to guarantee a $6/7$-approximation to the original MMS, and $\lfloor n/3\rfloor$ copies suffice for a $4/5$-approximation. Both results improve upon the best known approximation guarantees for additive valuations in the absence of copies.
@@ 1 Introduction
Fair Division goes back to antiquity, (Genesis 13), where 2 agents (Abraham and Lot) used cut and choose to split a contested resource. For divisible items and 2 agents, cut and choose ensures that (a) both agents get at least half the good, when viewed subjectively based on the agent’s own valuation — this is called a proportional allocation, and (b) no agent would prefer to swap with the other — this is called an envy free allocation. See [17, 18]. Extensions to $n>2$ agents were studied by Banach, Knaster and Steinhaus, [42], where proportionality means that every agent gets at least $1/n$ fraction of her value for the entire good. In particular, they show that proportionality is always achievable, for any number of agents, in the context of divisible cake cutting. (In general, for more than 2 agents, proportionality does not imply envy freeness.) Unfortunately, proportionality is not possible, in general, for indivisible items.
The maximin share is a relaxation of proportionality for indivisible items. Given $n$ agents and a set $M$ of indivisible items, the maximin share of an agent is the maximum value (over all partitions of $M$ into $n$ bundles), of the minimum valued bundle in the partition.
The maximin share (hereinafter denoted MMS), was studied by [32, 21]. Budish [21] defined the term and showed that if the number of agents goes to infinity and there are many copies of every item (also going to infinity) then one can approximate the maximin share to within a small error (going to zero). This was particularly relevant in the context of registration to classes, his motivating example.
Again, unfortunately, even if restricted to additive valuations, a maximin allocation need not exist [34, 27]. Albeit, there are several special cases where achieving the maximin share is in fact achievable, see [16].
Relaxations of MMS
@@
Given that achieving the maximin share is impossible, several relaxations have been studied. Implicitly introduced in [21], a $1$-out-of-$d$ maximin share is the maximum value each agent can guarantee while partitioning the set $M$ into $d\geq n$ bundles (instead of $n$ bundles in the original MMS definition), and receiving a minimum valued bundle.
Clearly, the larger $d$ is, the easier it is to achieve the benchmark. The question then becomes, what is the smallest value of $d\geq n$ such that there is an allocation for $n$ agents so that each agent achieves the $1$-out-of-$d$ maximin share. This is also referred to as the ordinal approximation. In [33] the authors show that $1$-out-of-$\lceil 3n/2\rceil$ MMS allocations exist for any instance with additive valuations, later improved to $1$-out-of-$4\lceil n/3\rceil$ MMS allocations in [4], the current state of the art. For all we know, even $1$-out-of-$(n+1)$ MMS allocations may exist.
@@
Another relaxation of MMS, also called approximate-MMS, is that of $\alpha$-MMS, introduced by Procaccia and Wang [34]. An $\alpha$-MMS allocation is one where every agent gets at least an $\alpha$ fraction of the maximin share. Procaccia and Wang [34] showed that $\alpha=2/3$ is attainable for additive valuations. This was subsequently improved in a sequence of papers, culminating in [1] which achieved a $(\frac{3}{4}+\frac{3}{3836})$-MMS allocation, the best known bound thus far (Akrami and Garg).
1.1 Our Contributions
Our main contribution is to introduce relaxations to the MMS by allowing resource augmentation, (i.e., allowing extra copies of items) and to obtain both positive and negative results in different settings. This relaxation was implicitly suggested by Budish in [21], in the context of course allocation (where the items are seats in a classroom), and it may also be applicable to other settings in which adding copies of an item is possible (albeit with some restrictions), e.g., assigning private tutors to students, caregivers to elderly individuals, or distributing any form of hourly-based service. We ask:
Main question:
@@ 1.1 Our Contributions
How many items should be duplicated so as to guarantee an allocation in which every agent’s value is at least the agent’s (original) MMS?
We consider both the total number of copies and the maximum number of copies of an item. We distinguish between distinct copies (only one extra copy of an item allowed), and when more copies of a single item are allowed. Much of this paper is focused on distinct copies — up to one extra copy of an item. Our results are the following:
General Monotone Valuations
A tight bound ($\approx m/e$) on the number of copies needed to achieve MMS with copies, for arbitrary monotone valuations. More precisely we show that
  • For all $n$, $m$, there exist $n$ monotone valuations over $m$ items such that every MMS allocation with copies requires at least $\left(\frac{n-1}{n}\right)^{n}\cdot m$ copies. See Theorem 1. The number of copies of an individual good can be as high as $n$ (in an instance where $m$ is sufficiently large).
@@
  • Given $n$ monotone valuations $\{v_{i}\}_{i\in[n]}$ and a set $[m]$ of goods, one can find an MMS allocation with copies, while making at most $\left(\frac{n-1}{n}\right)^{n}\cdot m$ extra copies. See Proposition 1. Moreover, the number of copies of any individual good is no more than $O(\ln m/\ln\ln m)$. See Proposition 2. Note that this does not conflict with the lower bound of $n$ above as the lower bound instance has $m=n^{n}$.
  • We also show a connection between two different MMS relaxations, in particular we show that if there exists a 1-out-of-$(1+\alpha)n$ MMS allocation then there is a full MMS allocation with $\lfloor\alpha m\rfloor(1+o(1))$ distinct copies, see Lemma 5. While this reduction is for general valuations — we use it to obtain bounds for additive valuations.
Additive Valuations
  • We show that
    $\min\{n-2,\lfloor m/3\rfloor(1+o(1))\}$
    distinct copies suffice to guarantee a full MMS allocation, see Theorem 3. These results are indeed tight for 3 additive agents (namely, a single copy of a single good suffice), but are not known to be tight for 4 or more agents, raising a very perplexing open problem. As far as we know today, a single copy of a single good may suffice to achieve a full MMS allocation for any instance with additive valuations.
  • We also consider MMS problem for chores, where the question becomes how many chores need be removed to achieve an MMS allocation? We show that there is an MMS allocation for chores where at least $m-n+2$ of the chores will be allocated to agents, (i.e., $n-2$ chores removed), see Appendix A.
@@
We give new approximation guarantees for additive instances, which improve upon the best-known approximations without copies. We give points along a tradeoff between the guarantee of the $\alpha$-MMS approximation and the number of copies needed. We show how to achieve a $6/7$-MMS with $\lfloor n/2\rfloor$ distinct copies (Theorem 4), and how to achieve a $4/5$-MMS with $\lfloor n/3\rfloor$ distinct copies (Theorem 5).
Relation to 1-out-of-$d$ MMS.
As mentioned, we show 1-out-of-$(1+\alpha)n$ MMS implies that exact MMS can be achieved with $\alpha m(1+o(1))$ distinct copies (see Lemma 5). In the other direction, showing that $n/\beta$ copies are sufficient for achieving exact MMS, immediately implies that 1-out-of-$(1-\beta)n$ MMS allocations exist. This is because at least $(1-\beta)n$ agents do not receive any copies, and obtain their MMS value. In particular, for additive valuations, achieving MMS with less than $n/4$ copies improves upon the currently best-known bound for 1-out-of-$d$.
1.2 Our Techniques
Our upper bounds for general monotone valuations use the probabilistic method. This is used for bounding the total number of copies as well as for bounding the maximum number of copies of a single item.
We give new extended variants of existing techniques in the fair division literature to a setting with copies. Specifically:
  • We introduce bag filling with copies. Our bag filling techniques give full MMS allocations while duplicating some of the items. See Algorithm 2: $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$.
  • Valid reductions in general take an instance and reduce it to a new instance with less agents and items, while ensuring that the MMS of each remaining agent has not decreased. An $\alpha$-valid reduction is a valid reduction where the agents removed are allocated sets whose value is at least $\alpha$ times their MMS share, see Definition 6. We give new valid reductions and $\alpha$-valid reductions that use copies, see Section 5.2.
@@ 1.2 Our Techniques
  • While it was shown by [16] that considering agents with the same order of valuations over the items is without loss of generality when studying MMS, their proof does not carry over to the setting with copies. Nevertheless, using an augmenting path argument, we prove that for a restricted class of allocations, in particular those produced by our algorithms, such a reduction applies (see Section 5). Consequently, our algorithms for achieving approximate MMS assume, without loss of generality, that the instance is ordered. It is an intriguing open problem whether such a reduction applies in general.
@@
  • We combine bag filling and round robin approaches to achieve approximate MMS allocations with copies for ordered instances. See Algorithm 4: $\mathsf{BagFill\mbox{-}RoundRobin}$.
@@ 1.3 Open Problems
We believe that one of our major contributions is to suggest many open problems, the penultimate open problem being “does one extra copy of a good suffice to obtain an MMS allocation for additive valuations?”. If so, this would be an interesting analogy to the Bulow Klemperer result [22] where adding a single bidder suffices so that a simple auction (Vickrey) obtains as much revenue as the optimal auction (Myerson). Other open problems are:
  • Of course, making progress on any of our bounds on additive valuations are obvious open problems, even if not showing that one extra copy suffices for MMS.
  • Considering more general classes of valuations such as submodular valuations, gross substitutes, etc. An illustrative example is that of $k$-demand valuations where we show that $n-1$ distinct copies suffice to achieve an MMS allocation, see Appendix D. We remark that the tools used to produce an MMS allocation with copies for $k$-demand valuations are a slight variant of those used for additive valuations.
  • We reduce the problem of MMS with [distinct] copies to the $1$-out-of-$d$ problem (Lemma 5). We use this to derive bounds on the number of copies in the context of additive valuations, but any such $1$-out-of-$d$ result for arbitrary monotone valuations would similarly imply analogous results for general valuations. Are there such $1$-out-of-$d$ results?
  • Another natural direction is applying the resource augmentation paradigm to other fairness notions.

### v3

@@ Abstract
We introduce and formalize the notion of resource augmentation for maximin share (MMS) fairness for the allocation of indivisible goods. Given an instance with $n$ agents and $m$ goods, we ask how many copies of the goods should be added in order to guarantee that each agent receives at least their original MMS value, or a meaningful approximation thereof.
  • For general monotone valuations, we establish a tight bound: an exact MMS allocation can be guaranteed using at most $\Theta(m/e)$ total copies, and this bound is tight even for XOS valuations. We further show that it is unavoidable to duplicate some goods $\Omega(\ln m/\ln\ln m)$ times, and provide matching upper bounds.
  • For additive valuations, we show that at most $\min\{n-2,\left\lfloor\frac{m}{3}\right\rfloor(1+o(1))\}$ distinct copies suffice. This separates additive valuations from submodular valuations, for which we show that n1n-1 copies may be necessary.
  • We also study approximate MMS guarantees for additive valuations and establish new tradeoffs between the number of copies needed and the approximation guaratee. In particular, we prove that $\lfloor n/2\rfloor$ copies suffice to guarantee a $6/7$-approximation to the original MMS, and $\lfloor n/3\rfloor$ copies suffice for a $4/5$-approximation. Both results improve upon the best-known approximation guarantees for additive valuations in the absence of copies.
  • Finally, we relate MMS with copies to the relaxed notion of 1-out-of-$d$ MMS, showing that improvements in either framework translate directly to the other. In particular, we establish the first impossibility results for 1-out-of-$d$ MMS.
Our results highlight the power and limits of resource augmentation for achieving MMS fairness.
[…]
@@ 1 Introduction
Fair Division dates back to antiquity, (Genesis 13), where two agents (Abraham and Lot) used cut and choose [Brams_Taylor_95, Brams_Taylor_1996] to split a contested resource. A natural notion of fairness is proportionality, which requires that each of the $n$ participating agents receives at least $1/n$-fraction of her value for the entire set of resources [steinhaus1948problem]. Unfortunately, for indivisible goods an allocation that achieves proportionality simultaneously for all agents may not exist.
A prominent relaxation of proportionality is the maximin share (MMS) notion [Hill87, budish2011combinatorial]. Given $n$ agents and a set $M$ of indivisible items, the maximin share of an agent is defined as the maximum value she can guarantee herself by partitioning the set of items $M$ into $n$ bundles and then receiving the least valuable bundle. Alas, even when restricted to additive valuations, where the value of a bundle is the sum of the values of its items, an MMS allocation need not exist, even with only three agents [Kurokawa18, feige2021tight]. Nonetheless, MMS allocations do exist in several natural special cases; see [bouveret2016characterizing].
Relaxations of MMS.
@@
Given that achieving the maximin share is impossible, several relaxations of MMS have been studied. Implicitly introduced in [budish2011combinatorial], a 11-out-of-$d$ MMS is the maximum value each agent can guarantee while partitioning the set $M$ into $d>n$ bundles (instead of $n$ bundles in the original MMS definition), and receiving a minimum valued bundle. Clearly, the larger $d$ is, the easier it is to achieve the benchmark. The question then becomes, what is the smallest value of $d$ such that there is an allocation for $n$ agents so that each agent achieves the 11-out-of-$d$ maximin share. The current state of the art is that $1$-out-of-$4\lceil n/3\rceil$ MMS allocations exist [akrami2023improving].
@@
An $\alpha$-MMS allocation is one where every agent gets at least an $\alpha$ fraction of the maximin share. This MMS relaxation was defined by Procaccia and Wang [Kurokawa18]. They show that $\alpha=2/3$ is attainable for additive valuations. This was subsequently improved in a sequence of papers, culminating in the recent work of [huang2025fptas79], which achieved a $\frac{7}{9}$-MMS allocation, the best known bound thus far.
MMS via Resource Augmentation.
In this work we take a resource augmentation approach to overcome the potential absence of MMS allocation. Namely, duplication of some of the existing goods.
The idea of using copies of goods to achieve the MMS follows naturally from the work of Budish [budish2011combinatorial], that considered a fair allocation of courses to students. Budish introduced a mechanism with strong fairness guarantees, but with a (vanishing) “market clearing error”, i.e., the mechanism may allocate more than the number of available seats in a class.
Resource augmentation is a central tool in computer science for beyond worst-case analysis [sleator1985amortized, Roughgarden_2021]. A prominent example is machine scheduling, where resource augmentation has been employed by increasing machine speed, adding machines, providing extra memory, and more, thereby improving performance relative to the benchmark of the non-augmented settings, or circumventing known impossibility results [extraSpeedMahcines00, chekuri2004multi, Albers04].
A canonical example of resource augmentation in the economic literature comes from [bulowklemperer96] in the context of revenue maximization in auctions, where the augmented resources are additional bidders. They show that in a single-item auction with $n$ i.i.d. bidders, adding just one more bidder ensures that the revenue of a simple auction exceeds the optimal revenue attainable in the original $n$-bidder market, by a complex and arguably non-intuitive mechanism. This result inspired a sequence of follow up works [eden2017competition, liu2018competition] on what is known as “competition complexity”. In fact, the notion of 1-out-of-$d$ MMS is somewhat related to the competition complexity measure, where large $d$ reflects high competition, leading to lower (easier) benchmarks.
More recently, resource augmentation has been studied in problems at the intersection of computer science and economics, including selfish routing [roughgarden2002bad, roughgarden2021beyond], seeding in social networks [akbarpour2025just, akbarpour2018diffusion], facility location [caragiannis2016truthful], and prophet inequalities [Brustle0DEFV24, BrustleCDV24].
1.1 Our Contributions and Techniques
Our main contribution is the introduction of a formal framework for resource augmentation, via item duplication, for the maximin share (MMS) fairness notion in fair division. Within this framework, we present both positive and negative results for different classes of valuation functions, shedding light on the possibilities and limitations of resource augmentation for achieving MMS guarantees. Our main research question is the following:
@@ 1.1 Our Contributions and Techniques
How many items must be duplicated to guarantee the existence of an allocation in which every agent receives value of at least her (original) MMS?
In our analysis, we consider two measures of duplication: (i) the total number of additional copies, and (ii) the maximum number of copies of any single item. When allowing at most one extra copy per item, we refer to this setting as having distinct copies. Our results are presented below and partly summarized in Table 1.
Additive Submodular General (Monotone)
Upper Bound
$\min\{n-2,\left\lfloor\frac{m}{3}\right\rfloor(1+o(1))\}$
(Thm. 3.6)
$m/e$
$m/e$
(Prop. 3.3)
Lower Bound
11
[Kurokawa18]
n1n-1
(Prop. 3.14)
$m/e$ (even for XOS)
(Thm. 3.1)
Table 1: Bounds on the total number of copies required to achieve exact MMS. Results for additive valuations hold even for distinct copies.
General Monotone Valuations.
We establish the following bound on the number of copies sufficient to guarantee an MMS allocation for general monotone valuations. Interestingly, this bound is tight, even for XOS valuations.
Main Theorem 1 (General monotone valuations, Theorem 3.1 and Proposition 3.3).
@@
For any set of $n$ agents with monotone valuations $\{v_{i}\}_{i\in[n]}$, and any set of $m$ goods, one can find an MMS allocation with copies, while making at most $\left(\frac{n-1}{n}\right)^{n}\cdot m$ extra copies. Moreover, this bound is tight, even for XOS valuations.
Our upper bounds are established via the probabilistic method, for both the total number of copies and per-item number of copies. Our lower bound is based on a careful construction of XOS valuations in which each agent $i$ has only $n$ designated bundles $P^{i}_{1},\dots,P^{i}_{n}$, with the property that any bundle guaranteeing her MMS value must contain one of these $P^{i}_{j}$ as a subset. Moreover, the intersection between any $P^{i}_{j}$ and $P^{i^{\prime}}_{j^{\prime}}$ for $i\neq i^{\prime}$ is large, which in turn forces the use of a large number of copies.
In addition to bounding the total number of copies, we show that the number of copies of any individual good is at most $O(\ln m/\ln\ln m)$, and that this bound is tight; see Theorem 3.4 and Lemma 3.5.111The theorem also shows that the number of copies of an individual good can be as high as $n$ (in an instance where $m$ is sufficiently large). Note that this result does not conflict with the above-mentioned lower bound of $n$, since the lower bound instance has $m=n^{n}$. Theorem 3.4 further asserts that, given MMS partitions for the agents, one can efficiently find—with high probability—an MMS allocation that satisfies the stated bounds on the extra copies.
Additive Valuations.
For additive valuations, all of our guarantees are achieved using distinct copies, namely, with at most one extra copy of each item. Our main result for additive valuations is as follows:
Main Theorem 2 (Additive valuations, Theorem 3.6).
For any instance with additive valuations, at most $\min\{n-2,\lfloor m/3\rfloor(1+o(1))\}$ distinct copies suffice to guarantee an exact MMS allocation.
We achieve our upper bound by combining a novel bag-filling procedure which introduces additional copies together with a preliminary matching stage. Technically, this involves considering a set of minimal size which violates Hall’s condition. This yields a tight bound for 3 additive agents (namely, a single copy of a single good is required to achieve MMS). Importantly, this bound is not tight for 4 or more agents, raising a very perplexing open problem. As far as we know today, a single copy of a single good may suffice to achieve an exact MMS allocation for any instance with additive valuations.
Interestingly, we show a separation between additive and submodular valuations, where n1n-1 copies are required to guarantee an MMS allocation.
Proposition (Lower bound for submodular valuations, Proposition 3.14).
For submodular valuations, any allocation with copies which achieves MMS requires at least n1n-1 total copies.
@@
We then turn to $\alpha$-MMS for additive valuations, where we obtain new approximation guarantees that improve upon the best-known results without copies. We identify points along a tradeoff curve between the achievable $\alpha$-MMS guarantee and the number of copies required. Our main results are stated in the following theorem.
Main Theorem 3 ($\alpha$-MMS for additive valuations, Theorem 4.2 and 4.3).
For any instance with additive valuations, $\lfloor n/2\rfloor$ copies suffice to guarantee a $6/7$-approximation to the original MMS, and $\lfloor n/3\rfloor$ copies suffice for a $4/5$-approximation.
Two common techniques in the study of MMS allocations are $\alpha$-valid reductions, and considering ordered instances.
An $\alpha$-valid reduction modifies an instance by removing a set of agents, each with an associated bundle that achieves an $\alpha$-fraction of her MMS value, while ensuring that the MMS values of the remaining agents do not decrease. Our results use novel $\alpha$-valid reductions for setting with copies, see Section 4.2.
@@
An ordered instance is one where all agents have identical ranking over the items. It was shown in [bouveret2016characterizing] that, one may assume, without loss of generality, that instances are ordered. Unfortunately, their argument does not carry over to settings with copies. Nevertheless, using an augmenting path argument, we prove that for a restricted class of allocations, in particular those produced by our algorithms, such a reduction applies (see Section 4.1). Consequently, our algorithms for achieving approximate MMS assume, without loss of generality, ordered instances. It is an intriguing open problem whether such a reduction applies in general.
@@
Combining our new $\alpha$-valid reductions with the ordered instances assumption, our algorithm further uses bag filling and round robin approaches to achieve $\alpha$-MMS allocations with copies; see Algorithm 4.
Relation to other relaxations of MMS.
We next relate our notion of MMS with copies to other relaxations of MMS. We start with a relation to the notion of 1-out-of-$d$ MMS, which holds for any class of valuations. We show that 1-out-of-$(1+\beta)n$ MMS implies that exact MMS can be achieved with $\beta m(1+o(1))$ distinct copies (see Lemma 3.10). In the other direction, one can show that, if $n\beta$ copies are sufficient for the existence of exact MMS, then 1-out-of-$\frac{n}{1-\beta}$ MMS allocations exist. This is because at least $(1-\beta)n$ agents do not receive any copies, and obtain their MMS values.
An interesting consequence for additive valuations is that achieving MMS with less than $n/4$ copies improves upon the currently best-known bound for 1-out-of-$d$, so such an improvement may be non-trivial.
Moreover, the example in Theorem 3.1 which shows the lower bound of $(\frac{n-1}{n})^{n}\cdot m$ on the number of extra copies, also gives the following intriguing impossibility result for the existence of 11-out-of-$d$ MMS allocations for XOS valuations (whose proof is deferred to Appendix A). Notably, this is the first impossibility result for 1-out-of-$d$ MMS.
Additionally, for subadditive valuations, [feige2025multi] showed that an $\alpha$-MMS allocation where each good is copied at most $k$ times (so the total number of copies is at most $nk$), implies the existence of a $\frac{\alpha}{4k}$-MMS allocation with no copies. In particular, proving the existence of such an allocation for a constant $\alpha$ and $k$, will improve upon the best known $\Omega(1/\log\log n)$-MMS for subadditive valuations given by [feige2025multi].
Main Theorem 4 (Impossibility of 1-out-of-$d$ for XOS valuations).
There exists no function $f:\mathbb{N}\rightarrow\mathbb{N}$ such that for all fair division instances $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ with $n$ agents and XOS valuations, a 11-out-of-$f(n)$ MMS allocation exists.
This impossibility result stands in stark contrast to $\alpha$-MMS guarantees for XOS valuations. Indeed, it is shown in [feige2025fair] that a constant approximation (precisely, $4/17$) exists for XOS valuations.
@@ 1.2 Open Problems
Our framework suggests a number of intriguing open problems. We highlight several directions below.
  • The preeminent open problem is whether adding just one extra copy of a single good suffices to guarantee an MMS allocation for additive valuations. More broadly, a natural direction is to further narrow the existing gaps for additive valuations.
  • We reduce the problem of MMS with distinct copies to the 11-out-of-$d$ problem (Lemma 3.10). Consequently, any positive 11-out-of-$d$ result for a given valuation class would immediately imply an upper bound on the number of copies required to guarantee full MMS for that class. While we establish an impossibility result for XOS valuations (Theorem 4), it remains open whether any positive or negative results exist for narrower valuation classes, such as submodular or gross-substitutes.
  • Another natural and largely unexplored direction is the design of truthful mechanisms that achieve MMS (or approximate MMS) allocations under resource augmentation. In particular, can one design incentive-compatible mechanisms that guarantee MMS allocations with copies?
  • In standard resource-augmentation frameworks, performance is measured relative to the original instance. An alternative direction is to study the effect of adding copies and evaluating fairness with respect to the new market.
  • Our algorithms for additive valuations duplicate at most one copy per item and aim to minimize the total number of copies. A different approach is to allow multiple copies of a single item. Even if up to $n$ copies of one item are permitted, determining the minimum total number of copies required to guarantee MMS remains open.
  • More generally, it would be interesting to explore the resource-augmentation paradigm in the context of other fairness notions, such as EF/EFX.